Some qualitative aspects of a nonlinear heat equation with variable coefficients

Massimo Borelli, Trieste, Italy

Keywords: Nonlinear fast diffusion, variable coefficients, instantaneous shrinking AMS Subject classification: 35K65

Abstract: Using the methods exposed by Kersner and Nicolosi, one extends their results on the instantaneous shrinking of the support of the non negative solutions to a class of singular parabolic equation with fast diffusion and strong absorption terms.